10 g of a mixture of sodium chloride and anhydrous sodium carbonate were made up to 1 dm3 of aqueous solution. 25cm3 of this solution required 20 cm3 of 0.2 M hydrochloric acid for neutralization.What was the mass of sodium chloride in the mixture
1
Expert's answer
2020-08-02T15:48:26-0400

25 cm3 of this solution required 20 cm3 of 0.2 M ( = 0.004 mol) hydrochloric acid for neutralization. Therefore, 1 dm3 of solution required 1000 * 0.004 / 25 = 0.16 mol of hydrochloric acid. It means that there is 0.16 / 2 = 0.08 mol of Na2CO3 in 10 g of a mixture. It is equivalent to 0.08 * 106 = 8.48 g of compound. This means that the remaining NaCl weighs 10 - 8.48 = 1.52 g.